OpenStudy (anonymous):

In right ∆ABC, m B = 90°, m C = 40°, and BC = 10. What are the other two side lengths of the triangle?

OpenStudy (anonymous):

help me its urgent

OpenStudy (anonymous):

Based on your given details the drawing would be:  To get the hypotenuse of the triangle which is AC you need to do the cos function (Make sure your calculator is in Degree mode!) It would look like this: \[\cos40 = \frac{ AC }{ 10 }\] The other leg of the triangle (AB) would be the tan function or if you know both BC and AC you could do the Pythagorean Theorem: \[a^2 + b^2 = c^2\] If you do tangent then your answer would be gotten like this: \[\tan 40 = \frac{ AB }{ 10 }\] Hope this helps!